Based on the Baseball camp example covered in the class, let's assume the segment size is 10,000, 1.5% of the the segment participate the camp, frequency is 1, variable cost per person is $5, TFC = $9,000, and price per person is unknown. Based on the assumption provided above, if the program wants to make $1500 profit, how much should they charge to each participant?

89?

75?

56?

91?
Answer #1

No. of people participation in the camp = 10000*1.5% = 150

Let, price to be charged = X

Then,

Net profit = Total Revenue – variable cost – fixed cost

1500 = X*150 – 5*150 – 9000 = (X-5)*150 – 9000

X = (1500+9000)/150 + 5

X = $75

So, $75 should be charged to each participant to earn $1500 profit.